OpenStudy (anonymous):

What is the quotient in simplified form? State any restrictions on the variable.

OpenStudy (anonymous):

\[\frac{ x^2 - 16 }{ x^2 + 5x + 6 } \div \frac{ x^2 + 5x + 4 }{ x^2 - 2x - 8 }\]

OpenStudy (anonymous):

Multiple choice answers: \[ \frac{ (x-4^2) }{ (x+3)(x+1) }\] x =/ -3 -1 \[\frac{ (x+4)^2(x+1) }{ (x+2)^2 (x+3) }\] x =/ -3, -2 4

OpenStudy (anonymous):

\[\frac{ (x-4)^2 }{ (x+3)(x+1) }\] x=/ -4 -3 -2 -1 4

OpenStudy (anonymous):

\[\frac{ 1 }{ (x+3(x+1) }\] x =/ -4 -3 -2 -1 4

hartnn (hartnn):

ok, first use this \(\dfrac{a}{b} \div \dfrac{c}{d} = \dfrac{a}{b} \times \dfrac{d}{c}\)

hartnn (hartnn):

so, \(\frac{ x^2 - 16 }{ x^2 + 5x + 6 } \div \frac{ x^2 + 5x + 4 }{ x^2 - 2x - 8 }=.. ?\)

hartnn (hartnn):

not required...just observe that if we change division sign to multiplication sign, the 2nd fraction flips...

OpenStudy (anonymous):

ooooh yup i c that

hartnn (hartnn):

flips means the numerator becomes denom., and denom. becomes num.

hartnn (hartnn):

\(\frac{ x^2 - 16 }{ x^2 + 5x + 6 } \div \frac{ x^2 + 5x + 4 }{ x^2 - 2x - 8 }=.. ?\)

OpenStudy (anonymous):

I have to find out what it equals?

OpenStudy (anonymous):

how would i start doing that?

hartnn (hartnn):

i meant , from the above explanation, \(\large \frac{ x^2 - 16 }{ x^2 + 5x + 6 } \div \frac{ x^2 + 5x + 4 }{ x^2 - 2x - 8 }=\frac{ x^2 - 16 }{ x^2 + 5x + 6 } \times \frac{ x^2 - 2x - 8}{ x^2 + 5x + 4 }\) got this ?

OpenStudy (anonymous):

yup got it

hartnn (hartnn):

now do you also need help factoring those 4 expressions ?

OpenStudy (anonymous):

oh yeasss

hartnn (hartnn):

or can you factor by yourself ?

hartnn (hartnn):

\(x^2-16= x^2-4^2=(x+4)(x-4) \\as, a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

As a wild guess I think its A

OpenStudy (anonymous):

its C. -4 -3 -2 -1 4